# Selmer stability in families of congruent Galois representations

### v1

@@ 3. Main results
In this section, I fix a prime $p\geq 5$, a $p$-adic field $\mathcal{K}$ with valuation ring $\mathcal{O}$, uniformizer $\varpi$ and residue field $\kappa$. I also fix a surjective and odd Galois representation
$\bar{\rho}:\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{GL}_{2}(\kappa)$
@@
Let $\Omega_{\bar{\rho}}$ be the set of prime numbers such that
@@
Let $\ell\in\Omega_{\bar{\rho}}$ and $\varrho:\operatorname{G}_{\ell}\rightarrow\operatorname{GL}_{2}(\mathcal{O})$ be a continuous representation for which $\bar{\varrho}=\bar{\rho}_{f}$. Assume that $\varrho$ is ramified at $\ell$. Then, up to conjugation by a matrix in $\operatorname{GL}_{2}(\mathcal{O})$,
$\varrho(\sigma_{\ell})=\left({\begin{array}[]{cc}\ell x&\\ &x\\ \end{array}}\right)\text{ and }\varrho(\tau_{\ell})=\left({\begin{array}[]{cc}1&y\\ &1\\ \end{array}}\right),$
where $x\equiv-1\mod{\varpi}$ and $y\equiv 0\mod{\varpi}$.
Proof.
For a proof of this result, see [Ray24, Lemma 4.4]. ∎
Lemma 3.3.
@@
By Lemma 3.2, there exist $x\equiv-1\pmod{\varpi}$ and $y\equiv 0\pmod{\varpi}$ such that
@@
as a $\mathrm{G}_{\ell}$-module. The element $\sigma_{\ell}$ acts on $\kappa\cdot e_{2}$ via multiplication by $x\bmod\varpi$, which is $-1$ by assumption. Since $p\neq 2$, this action is nontrivial, and so the invariants under $\mathrm{G}_{\ell}/\mathrm{I}_{\ell}$ vanish:
@@
Let $g$ be a $p$-ordinary modular form of weight $2$ such that $\bar{\rho}_{g}\simeq\bar{\rho}_{f}\simeq\bar{\rho}$. Assume that $p\nmid N_{g}$. Let $\ell$ be a prime dividing $N_{\bar{\rho}}$ such that ±$\ell\not\equiv\pm 1\pmod{p}$. Then
@@
Since $\ell\mid N_{\bar{\rho}}$, the representation $\bar{\rho}$ is ramified at $\ell$, and in particular,
$H^{0}(\mathrm{I}_{\ell},A_{g}[\varpi])\neq 0.$
Thus, there exists a basis in which the restriction $\bar{\rho}|_{\mathrm{I}_{\ell}}$ takes the form
@@
so that $\bar{\rho}$ factors through the maximal pro-$p$ tame inertia quotient of $\mathrm{G}_{\ell}$, generated by $\sigma_{\ell}$ and $\tau_{\ell}$ with the relation $\sigma_{\ell}\tau_{\ell}\sigma_{\ell}^{-1}=\tau_{\ell}^{\ell}$. Here, $\sigma_{\ell}$ is a lift of Frobenius and $\tau_{\ell}$ is a generator of the tame inertia. Suppose $\bar{\rho}(\tau_{\ell})=\begin{pmatrix}1&y\\ 0&1\end{pmatrix}$ with $y\neq 0$, and write $\bar{\rho}(\sigma_{\ell})=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$. Using the relation above, I compute
@@
Equating these matrices, I obtain $c=0$ and $a=d\ell$. Hence,
$\bar{\rho}(\sigma_{\ell})=\begin{pmatrix}d\ell&b\\ 0&d\end{pmatrix}.$
Since $\det\bar{\rho}(\sigma_{\ell})=\ell$, I conclude that
$d^{2}\ell=\ell\quad\Rightarrow\quad d=\pm 1.$
Assume, for contradiction, that $\beta_{\ell}(A_{g})>0$. Then the same reasoning as in the proof of Lemma 3.3 implies that the image of $\sigma_{\ell}$ in $\mathrm{G}_{\ell}/\mathrm{I}_{\ell}$ acts trivially on the unique Jordan–Hölder constituent of $H^{0}(\mathrm{I}_{\ell},A_{g}[\varpi])$. That is, $d\ell\equiv 1\pmod{\varpi}$.
Since $d=\pm 1$, this would imply
$\ell\equiv d^{-1}\equiv\pm 1\pmod{p},$
contradicting the assumption that ±$\ell\not\equiv\pm 1\pmod{p}$. Therefore, $\beta_{\ell}(A_{g})=0$. The vanishing of $\beta_{\ell}(A_{f})$ follows by the same argument. ∎
@@
    No prime ±$\ell\equiv\pm 1\pmod{p}$ divides $N_{\bar{\rho}}$.
@@
This will imply the result, since $\bar{\rho}_{f}\simeq\bar{\rho}_{g}$.
@@
$0\to\mathrm{Sel}_{\mathrm{Gr}}(A_{h}[\varpi]/\mathbb{Q})\to\mathrm{Sel}_{\mathrm{Gr}}(A_{h}/\mathbb{Q})[\varpi]\to\bigoplus_{\ell}\operatorname{ker}h_{\ell},$
I have the inequality
@@
By Lemma 2.8, I have $\beta_{\ell}(A_{h})=0$ for $\ell=p$ and for all $\ell\nmid N_{h}$. For the remaining primes $\ell\mid N_{h}$, there are two cases:
@@
  • If $\ell\mid N_{\bar{\rho}}$, then ±$\ell\not\equiv\pm 1\pmod{p}$ by assumption. Lemma 3.4 applies, and again I conclude $\beta_{\ell}(A_{h})=0$.
Hence, I obtain the desired equality:
@@
Before proving our main theorem, I recall the following well known result of Serre.
@@
Let $\Omega:=\Omega_{\bar{\rho}}$, and consider any $M\in M_{\Omega}(Y)$. Set $N:=MN_{\bar{\rho}}$. Then $N\in\mathcal{S}(\bar{\rho})$, and by Theorem 2.4, there exists a Hecke eigencuspform $g$ such that $\bar{\rho}_{g}\simeq\bar{\rho}_{f}$ and $N=N_{g}$. By Proposition 3.5, I have
@@
Applying Theorem 3.6, I obtain

### v2

@@ 3. Main results
In this section, we fix a prime $p\geq 5$, a $p$-adic field $\mathcal{K}$ with valuation ring $\mathcal{O}$, uniformizer $\varpi$ and residue field $\kappa$. We also fix a surjective and odd Galois representation
$\bar{\rho}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{GL}_{2}(\kappa)$
@@
Let $\Omega_{\bar{\rho}}$ be the set of prime numbers $\ell$ such that
@@
It follows from [RAY24, Lemma 4.4] that $\rho_{g}$ factors through the maximal pro-$p$ tame inertia quotient of $\mathrm{G}_{\ell}$, generated by $\sigma_{\ell}$ and $\tau_{\ell}$ with the relation $\sigma_{\ell}\tau_{\ell}\sigma_{\ell}^{-1}=\tau_{\ell}^{\ell}$. Here, $\sigma_{\ell}$ is a lift of Frobenius and $\tau_{\ell}$ is a generator of the tame inertia. Further, there exist $x\equiv-1\pmod{\varpi}$ and $y\equiv 0\pmod{\varpi}$ such that
@@
as a $\mathrm{G}_{\ell}$-module. The element $\sigma_{\ell}$ acts on $\kappa\cdot e_{2}$ via multiplication by $x\bmod\varpi$, which is 1-1 by assumption. Since $p\neq 2$, this action is nontrivial, and so the invariants under $\mathrm{G}_{\ell}/\mathrm{I}_{\ell}$ vanish:
@@
Let $g$ be a $p$-ordinary modular form of weight $2$ such that $\bar{\rho}_{g}\simeq\bar{\rho}_{f}\simeq\bar{\rho}$. Assume that $p\nmid N_{g}$. Let $\ell$ be a prime dividing $N_{\bar{\rho}}$ such that $\ell\not\equiv\pm 1\pmod{p}$. Then
@@
Let $h$ denote $f$ or $g$. Assume by way of contradiction that $\beta_{\ell}(A_{h})>0$. By Part (1) of Lemma 2.7 we have $H^{0}\left(I_{\ell},A_{h}\right)\neq 0$, hence $H^{0}\left(I_{\ell},A_{h}[\varpi]\right)\neq 0$. Since $\det\bar{\rho}$ is the mod $p$ cyclotomic character and $\ell\neq p$, we find that $\operatorname{det}(\bar{\rho}_{|\operatorname{I}_{\ell}})=1$. Consequently, there exists a basis $(\bar{e_{1}},\bar{e_{2}})$ with respect to which the restriction $\bar{\rho}|_{\mathrm{I}_{\ell}}$ takes the form
@@
so that $\bar{\rho}$ factors through the maximal pro-$p$ tame inertia quotient of $\mathrm{G}_{\ell}$, generated by $\sigma_{\ell}$ and $\tau_{\ell}$ (as in the proof of Lemma 3.2). Since $\ell\mid N_{\bar{\rho}}$, the representation $\bar{\rho}$ is ramified at $\ell$. Therefore there exists $y\neq 0$ for which $\bar{\rho}(\tau_{\ell})=\begin{pmatrix}1&y\\ 0&1\end{pmatrix}$. We also write $\bar{\rho}(\sigma_{\ell})=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$. From the relation $\sigma_{\ell}\tau_{\ell}\sigma_{\ell}^{-1}=\tau_{\ell}^{\ell}$, we find that
@@
Equating these matrices, we obtain $c=0$ and $a=d\ell$. Hence,
(3.2) $\bar{\rho}(\sigma_{\ell})=\begin{pmatrix}d\ell&b\\ 0&d\end{pmatrix}$.
Note that $\det\bar{\rho}(\sigma_{\ell})=\bar{\chi}(\sigma_{\ell})=\ell$, and thus we conclude that $d^{2}\ell=\ell$, and consequently, $d=\pm 1$.
Since $\bar{\rho}\left(\tau_{\ell}\right)-1=\left(\begin{array}[]{ll}0&y\\ 0&0\end{array}\right)$, we see that $\rho\left(\tau_{\ell}\right)$ has Smith Normal Form $\left(\begin{array}[]{cc}1&0\\ 0&w\end{array}\right)$ for some $w\in(\varpi)$. If $w=0$ then $A_{h}^{I_{\ell}}/\varpi A_{h}^{I_{\ell}}=0$, contradicting $\beta_{\ell}\left(A_{h}\right)>0$. Thus $w\neq 0$, giving $A_{h}^{I_{\ell}}=\left(\varpi^{-N}\mathcal{O}/\mathcal{O}\right)\cdot\gamma$ for some $N>0$ and some $\gamma$ which is necessarily congruent to $\bar{e}_{1}$ modulo $\varpi$. Since $\ell\neq\pm 1(\bmod\varpi)$, we see from (3.2) that the action of $\sigma_{\ell}-1$ on $A_{h}^{I_{\ell}}/\varpi A_{h}^{I_{\ell}}$ is invertible, hence $\left(A_{h}^{I_{\ell}}/\varpi A_{h}^{I_{\ell}}\right)^{G_{\ell}/I_{\ell}}=0$. Part (1) of Lemma 2.7 gives $\beta_{\ell}\left(A_{h}\right)=0$. This is a contradiction and the proof is complete. ∎
@@
    No prime $\ell\equiv\pm 1\pmod{p}$ divides $N_{\bar{\rho}}$.
@@
(3.3) $0\to\mathrm{Sel}_{\mathrm{Gr}}(A_{h}[\varpi]/\mathbb{Q})\to\mathrm{Sel}_{\mathrm{Gr}}(A_{h}/\mathbb{Q})[\varpi]\to\bigoplus_{\ell}\operatorname{ker}h_{\ell},$
we have the inequality
@@
By Lemma 2.7, we have $\beta_{\ell}(A_{h})=0$ for $\ell=p$ and for all $\ell\nmid N_{h}$. For the remaining primes $\ell\mid N_{h}$, there are two cases:
@@
  • If $\ell\mid N_{\bar{\rho}}$, then $\ell\not\equiv\pm 1\pmod{p}$ by assumption. Lemma 3.3 applies, and again we conclude $\beta_{\ell}(A_{h})=0$.
Hence, we obtain the desired equality:
(3.4) $\dim_{\kappa}\mathrm{Sel}_{\mathrm{Gr}}(A_{h}/\mathbb{Q})[\varpi]=\dim_{\kappa}\mathrm{Sel}_{\mathrm{Gr}}(A_{h}[\varpi]/\mathbb{Q}).$
The inertia group at $p$ acts on $A_{h}[\varpi]$ via $\left({\begin{array}[]{cc}\bar{\chi}&\ast\\ 0&1\\ \end{array}}\right)$ and $A_{h}[\varpi]^{+}$ is the $1$-dimensional submodule on which $\operatorname{I}_{p}$ acts via the mod-$p$ cyclotomic character and $A_{h}[\varpi]^{-}$ is the unique quotient on which $\operatorname{I}_{p}$ acts trivially. Thus $A_{h}[\varpi]$ determines the $1$-dimensional modules $A_{h}[\varpi]^{\pm}$. Since $\bar{\rho}_{f}\simeq\bar{\rho}_{g}$, we deduce that
$\dim_{\kappa}\mathrm{Sel}_{\mathrm{Gr}}(A_{f}[\varpi]/\mathbb{Q})=\dim_{\kappa}\mathrm{Sel}_{\mathrm{Gr}}(A_{g}[\varpi]/\mathbb{Q}).$
From (3.4) it follows that
@@
Before proving our main theorem, we recall the following well known result of Serre.
@@
Let $\Omega\mathrel{\mathop{\ordinarycolon}}=\Omega_{\bar{\rho}}$, and consider any $M\in M_{\Omega}(Y)$. Set $N\mathrel{\mathop{\ordinarycolon}}=MN_{\bar{\rho}}$. Then $N\in\mathcal{S}(\bar{\rho})$, and by Theorem 2.4, there exists a Hecke eigencuspform $g$ such that $\bar{\rho}_{g}\simeq\bar{\rho}_{f}$ and $N=N_{g}$. By Proposition 3.4, we have
@@
Applying Theorem 3.5, we obtain